OpenStudy (anonymous):

Hey! Do you think you could create some example problems for me and walk me through it so I learn how to solve questions like this?

OpenStudy (anonymous):

Factor completely 4x2 - 81 Here is the question could you make something like it so i can learn how to solve it?

OpenStudy (anonymous):

raised to the second power

OpenStudy (amistre64):

interactmath.com tons of practice problems, free enter, pick a textbook on your subject, and practice to your hearts content

Nnesha (nnesha):

\[\huge\rm 9x^2 - 64 \]

OpenStudy (amistre64):

if your just looking for a difference of squares .. a^2 x^2 - b^2 = (ax+b)(ax-b) its a formula that needs no real practice.

OpenStudy (anonymous):

So from this equation 4x2 - 81 what would go into the "ax" part of the formula?

OpenStudy (amistre64):

sqrt(4x^2)

OpenStudy (anonymous):

So would you just divide everything by two?

OpenStudy (amistre64):

no, youd just take the sqrt of the first term ... as indicated

OpenStudy (anonymous):

Because wouldnt it be (2x + 9)

OpenStudy (anonymous):

So take the square root of each number?

OpenStudy (amistre64):

yes; (2x+9) (2x-9)

OpenStudy (anonymous):

So my next question is how do you know if you need the addition sign or the minus sign?

OpenStudy (amistre64):

in which part are you refering? a^2x^2 - b^2 or the factorization? (ax+b)(ax-b)

OpenStudy (anonymous):

the factorization.. Like why is it (2x+9) (2x-9) rather than two addition signs or two minus signs?

OpenStudy (amistre64):

foil it out ...

OpenStudy (anonymous):

How would you do that? Because the equation 4x^2 - 81 only has a minus sign?

OpenStudy (amistre64):

foil ... it .... they teach you how to multiply binomials before they want you to factor them right?

OpenStudy (amistre64):

the best way to see it is to actually work the process ... multiply (2x+9) and (2x-9) see what happens along the way

OpenStudy (anonymous):

They did i under stand there is such thing as the foil method like first outer inner and last but i dont really remember how to do it..

OpenStudy (amistre64):

great, memorization of a memory trick is only as good as memorizing sooo many tricks ... then just use distribution

OpenStudy (amistre64):

(2x + 9) (2x - 9) distribute it 2x(2x - 9) + 9(2x - 9) and distrbute again 4x^2 - 18x + 18x - 81 what happens to the middle terms?

OpenStudy (anonymous):

Youd combine them so 36x^2?

OpenStudy (amistre64):

look at it again ... or better yet, you work it out instead of trying to follow it along ... there is something real obvious you are glancing over in this.

OpenStudy (anonymous):

would you just take it out?

OpenStudy (anonymous):

Im not really sure to be honest

OpenStudy (amistre64):

why take it out? -18x + 18x what happens to them?

OpenStudy (anonymous):

Wouldnt they cancel out because its a negative 18.. then you would add 18 back to it? so would it just be x^2?

OpenStudy (amistre64):

\[(2x+9)(2x-9)=4x^2 - 18x + 18x - 81\] \[4x^2 \underbrace{\color{red}{- 18x + 18x}}_{=0} - 81\implies 4x^2-81\]

OpenStudy (amistre64):

does this only happen in this single case? or for all cases?

OpenStudy (anonymous):

Ohh so they "x" would also cancel out?

OpenStudy (amistre64):

yep: -18(4) + 18(4) = 0 just as it would for any other value of x

OpenStudy (anonymous):

It would always equal 0 if its a negative number with a positive number

OpenStudy (anonymous):

But if they were both positive you would just add them together, correct?

OpenStudy (amistre64):

the simple consequence of doing: \[(ax+b)(ax-b)\implies a^2x^2-b^2\] its just becasue that is the way the math works out. if they are both positive, they have a different consequence of the distribution \[(ax+b)^2=(ax+b)(ax+b)\implies a^2x^2+b^2+2abx\]

OpenStudy (amistre64):

if they are both negative, the consequence is a -2abx on the end instead ....

OpenStudy (anonymous):

Ohh okay I gotcha :)

OpenStudy (anonymous):

Thank you Sir

OpenStudy (amistre64):

your welcome

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @yoyogators the factorization.. Like why is it (2x+9) (2x-9) rather than two addition signs or two minus signs? \(\color{blue}{\text{End of Quote}}\) if there is a negative sign in original question then answer always going to be ( # + # ) (#.. - ..#.) one plus and one minus sign :-)

Nnesha (nnesha):

4x^2 + 9 can you factor this one ??

OpenStudy (anonymous):

Would they both be positive?

Nnesha (nnesha):

for that one you can't USE *DIFFERENCE OF SQUARE METHOD* :-)

OpenStudy (anonymous):

So would it be similar to what he said above. Would it be like this? "a2x2+b2+2abx"

Nnesha (nnesha):

that equation is not factorable :-)

Nnesha (nnesha):

in order to use difference square method there should be negative sign and both therms should be perfect square root :-)